Find the value of k for which angle subtended a line segment joining the points P(-1,2) and Q(1,4)at the point R(k,0)is maximum?

Grade:12th Pass

1 Answers

ankitesh gupta
11 years ago

IF YOU WANT MAXIMUM ANGLE TO BE SUBTENDED THEN THE POINT R(k,0) HAS TO LIE ON LINE PQ SO THAT MAXIMUM ANGLE SUBTENDED IS 180 DEGREE THEREFORE EQUATION OF LINE PQ : Y-2=[(2-4)/(-1-1)]{X-(-1)}

 Y-2=X+1  ....................... R POINT SHOULD SATISFY R(k,0)

 0-2=K+1.......... THEREFORE K=-3..................ANSWER

 

 PLZZZZZZZZZZZ DO APPROVE THE ANSWER IF IT HELPED YOU BY CLICKING YES BUTTON